OpenStudy (anonymous):

what is the difference 79.73 -56.31

OpenStudy (anonymous):

@whiteandnerdy

OpenStudy (anonymous):

just subtract the 2 answers

OpenStudy (myth1111):

Just subtract the 2 numbers. :x

OpenStudy (blake57roger):

23.42

OpenStudy (anonymous):

thanks

OpenStudy (myth1111):

@blake57roger -_- You need to explain how to do it instead of giving the direct answer.

OpenStudy (anonymous):

79.73 -56.31 ------ 23.42 3-1=2 7-3=4 put the decimal 9-6=3 7-5=2

OpenStudy (anonymous):

0923-32